The marks (out of 100) obtained by a group of students in a science test are 85, 76, 90, 85, 39, 48, 56, 95, 81, and 75. Find the:
Mean marks obtained by the group.
 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

First we have to arrange the marks obtained by a group of students in a science test in an ascending order,
= 39, 48, 56, 75, 76, 81, 85, 85, 90, 95
Mean of Marks = (Sum of all marks obtained by the group of students)/
(Total number of marks)
= (39 + 48 + 56 + 75 + 76 + 81 + 85 + 85 + 90 + 95)/ 10 = 730/10
= 73

​​Thanks